The diffrence between two numbers is -38. Two times the smaller number minus five times the larger number is -217. What are the numbers
Answer by checkley71(8403)

DIFFERENCE MEANS SUBTRACTION. THE DIFFERENCE BETWEEN 5 & 10 IS 10-5=5.
OR THE DIFFERENCE BETWEEN 2 NUMBERS IS 5 & THE LARGEST NUMBER IS 10 THEN:
X-Y=5 & X=10
10-Y=5
-Y=5-10
-Y=-5
Y=5
-------------------------------------------------
X-Y=-38 OR X=Y-38
2X-5Y=-217
2(X-38)-5Y=-217
2X-76-5Y=-217
-3Y=-217+76
-3Y=-141
Y=-141/-3
Y=47 ANSWER.
X-47=-38
X=-38+47
X=9 ANSWER.
PROOF:
2*9=5*-47=-217
18-235=-217
-217=-217